Il. Analyze the hyperbola with center (1,4), focus (-2, 4), and vertex (0, 4). That is,t nd: a) the other vertex 2 4 b) the other two points c) the transverse axis I) the other focus (4,4) the equation of the graph -))2-04
Answer #1

The standard equation for a hyperbola with a horizontal transverse axis and center at (h,k) is (x-h)2/a2 – (y-k)2/b2 = 1. The distance between the vertices is 2a and the distance between the foci is 2c. Also, a2+b2 = c2 .

Here, the center, focus, and vertex all lie on the horizontal line y = 4( i.e. these points lie on a line parallel to the X-Axis ). Thus, the hyperbola has a horizontal transverse axis.

a).The vertices of the hyperbola are equidistant from the center so that the other vertex is at (2,4).

b). Which other 2 points?

c).The center, focus, and vertex all lie on the horizontal line y = 4 which is the transverse axis.

d). The focus is 3 units from the center, so c = 3. Hence, the other focus is (4,4).

e). The vertex is 1 unit from the center so that a = 1. Now, since a2+b2 = c2 , hence b2 = 9-1 = 8. Thus, the equation of the hyperbola is (x-1)2/1 - (y-4)2/8 = 1.